OpenStudy (anonymous):

HELP!!! GEOMETRY! Midpoint of ST is (5, -8). Find S if T is: a) T (0,4) b) T (5, -15) c) T (10, 18) d) T (-2, 8) e) T (1, 12)

OpenStudy (klimenkov):

\(x_m,y_m\) is between \((x_1,y_1)\) and \((x_2,y_2)\) if \(x_m=\frac{x_1+x_2}2\) and \(y_m=\frac{y_1+y_2}2\).

OpenStudy (anonymous):

can you show me A?

OpenStudy (anonymous):

-5, 16 ?? :))

OpenStudy (klimenkov):

We need to find the coordinates of S. Let them be \((S_x,S_y)\). Let the coordinates of T be \((T_x,T_y)\). They are (0,4). Let rhe coordinates of midpoint be \((ST_x,ST_y)\). Using the formula we get: \(ST_x=\frac{S_x+T_x}2=\frac{S_x+0}2=5\). So, \(S_x=10\). The same for \(S_y\): \(ST_y=\frac{S_y+T_y}2=\frac{S_y+4}2=-8\) So, \(S_y=-20\). Answer: (10,-20).

OpenStudy (klimenkov):

Did you understand?